OpenStudy (anonymous):

(2sinθcosθ-cosθ)/(1-sinθ+sin^2θ-cos⁡^2θ)

OpenStudy (anonymous):

Do you get \[cot(\theta)-2cos(\theta)?\] \[\frac{2sin(\theta)cos(\theta)-cos(\theta)}{1-sin(\theta)+sin^2(\theta)-sin^2(\theta)+1}\rightarrow-2cos(\theta)+cot(\theta)\]